# MIMIR: Masked Image Modeling for Mutual Information-based Adversarial Robustness

### v3

@@ 4.2 Experimental Results
100 $\rightarrow$ 300 epochs FT 71.00 46.10
Table 4: Comparison between HSIC (Gretton et al. 2005) and $I_{\alpha}$ (Yu, Yu, and Príncipe 2021) using ViT-T on CIFAR-10. Models are pre-trained 800 epochs and adversarially fine-tuned with 10-step PGD for 50 epochs.
@@
Table 5: Comparison between different pre-training settings. All models are pre-trained for 800 epochs and then fine-tuned with 10-step PGD for 50 epochs using ViT-T on CIFAR-10.
@@
MI Measure. In Section 3.4, we provide lower and upper bound (Eq. (11)) of $I(x+\delta,z)$. According to the two bounds, $I(x+\delta,z)$ is supposed to decrease while the autoencoder learns to reconstruct the natural image $x$. This motivates us to directly embed $I(x+\delta,z)$ as a minimizing learning objective. In this paper, we use $I_{\alpha}$ (Yu, Yu, and Príncipe 2021) and HSIC (Ma, Lewis, and Kleijn 2020) as estimators (detailed definitions in Appendix F). Table 4 demonstrates the performance of different values of $\lambda$. According to the results, we use HSIC with $\lambda=1e-05$ for all other experiments.
1-step is better than 10-step of AT in pre-training. We also show that MIMIR outperforms original MAE (He et al. 2022) and adv MAE with different PGD steps (to generate adversarial examples for training). MAE in Table 5 refers to using the original MAE for pre-training and then fine-tuning with 10-step PGD. The adv MAE refers to using adversarial examples without the MI $I(x+\delta,z)$ in loss. The adv MAE (10-steps) refers to using the 10-step PGD algorithm ($\epsilon=8,\alpha=2$) to generate adversarial examples at pre-training. The adv MAE provides higher accuracy than MAE, which supports our statement that using adversarial examples in Masked Image Modeling creates a more difficult reconstruction task. This more difficult task further improves the performance of downstream models (See also Table 7). We use the default learning rate (i.e., $45.0e-$ 4) of MAE, so there is a performance drop in experiments in Table 5 and 4 since AT prefers larger learning rates on CIFAR-10 as shown in Table 11 in the Appendix.
Stability. Due to the high computation cost, we cannot report standard deviation for all experiments. To show that our method MIMIR has low variances, we train ViT-S on CIFAR-10 three times with different random seeds, running pre-training 400 epochs and fine-tuning 50 epochs. The natural accuracy is 86.07 $\pm$ 0.16 %. The adversarial accuracy under PGD20 ($l_{\infty},\epsilon=8/255$) is 47.24 $\pm$ 0.12 %.
Table 6: Adversarial accuracy by adaptive attacks. The models are pre-trained 800 epochs by MIMIR and fine-tuned 100 epochs by PGD AT.
@@
Adaptive Attacks. We evaluate MIMIR against adaptive adversaries following common practices (Tramer et al. 2020). We provide two adaptive attacks specifically designed against MIMIR. First, we introduce the PGD Mutual Information attack (PGD-MI), which utilizes the MI $I(x+\delta,z)$ to generate adversarial examples, as $I(x+\delta,z)$ is used in MIMIR pre-training as a penalty in the loss. PGD-MI attacks the model by directly increasing the MI $I(x+\delta,z)$. Specifically, we add the MI loss into the PGD algorithm:
@@
where the value of $\lambda$ in MIMIR pre-training is available to adversaries.
Second, we introduce a PGD feature attack (PGD-fea) that directly attacks the feature extracted by ViT backbones following (Sabour et al. 2016). In particular, we attack the feature extractor from the backbones after the adversarial fine-tuning. The PGD-fea attack increases the Euclidean distance between features extracted from natural and adversarial examples. We implement it using the PGD algorithm:
@@
Both PGD-MI and PGD-fea are optimized for 100 steps to make sure the attacking algorithm converges. The perturbation budget is the same as the previous evaluation, i.e., $\epsilon=8/255$ for CIFAR-10 and Tiny-ImageNet, and $\epsilon=4/255$ for ImageNet-1K. Table 6 demonstrates the adaptive evaluation results for PGD-MI and PGD-fea attacks, showing that MIMIR is robust against adaptive attacks on benchmark datasets.
Table 7: Natural accuracy of MAE and MIMIR (800 epochs pre-training for both) that are fine-tuned on natural images.
Architecture Pre-train CIFAR-10 Tiny ImageNet
@@
Figure 3: The loss landscapes of MIMIR pre-trained models.
@@
Visualization of the Loss Landscape. To show that the robustness of MIMIR-trained models does not stem from gradient masking, we plot the loss landscape (Li et al. 2018) in Figure 3. The loss landscape is the visualization of the loss function as parameters change. The basic idea is to plot the loss around the optimal parameters. The landscapes of all models are smooth, i.e., the gradient at a certain point is clear and can also be easily estimated by local average gradients, which means the gradient is masked.
Fine-tuning with Natural Images. In Table 7, we show the results of MIMIR and the original MAE (He et al. 2022) with the same hyperparameter. We fine-tune 50 epochs for CIFAR-10 and Tiny-ImageNet, 100 epochs for ImageNet-1K. Surprisingly, MIMIR outperforms MAE when fine-tuning with natural data. We leave further exploration of this finding to future work.
5 Conclusions and Future Work
This paper provides a novel theoretical MI analysis on ViT adversarial training, showing that MI between the adversarial example and its latent representation in ViT-based autoencoders, i.e., $I(x+\delta,z)$, should be constrained by utilizing the MI bounds. Based on this finding, we propose MIMIR as a theoretically grounded pre-training method to improve adversarial robustness for ViTs. MIMIR uses adversarial examples as inputs and natural data as the reconstruction target. In this way, the information from the adversarial perturbations is decreased by the bottleneck, while the information from natural data is preserved while reconstructing the target. Our experimental results show that MIMIR substantially improves adversarial robustness compared to recent related works on various benchmark datasets.
Appendix A Datasets
We use three commonly used datasets to evaluate MIMIR: CIFAR-10 (Krizhevsky, Hinton et al. 2009), Tiny-ImageNet (Le and Yang 2015), and ImageNet-1K (Deng et al. 2009). CIFAR-10 (Krizhevsky, Hinton et al. 2009) comprises 50,000 images with size $3\times 32\times 32$ in 10 classes. ImageNet-1K (Deng et al. 2009) is the most commonly used dataset for the evaluation of ViTs and its variants, which is composed of more than 1.2 million high-resolution images in 1,000 classes. In our experiments, images from ImageNet-1K are resized to $3\times 224\times 224$. For completeness, we also include Tiny-ImageNet (Le and Yang 2015) as a medium size dataset between CIFAR-10 (Krizhevsky, Hinton et al. 2009) and ImageNet-1K (Deng et al. 2009). Tiny-ImageNet (Le and Yang 2015) contains 100,000 images with size $3\times 64\times 64$ in 200 classes.
Appendix B Decoder Hyperparameters
We use transformer blocks but fewer layers as the backbone of the decoder. For CIFAR-10, we use the patch size of 2, 4 for Tiny-ImageNet, and 16 for ImageNet-1K. Table 8 shows the hyperparameters of decoder architectures. For different ViT architectures, we use the transformer blocks of the respective architectures to build the encoder.
Table 8: Model architectures of the encoder and decoder.
Model Layers Hidden size MLP ratio Heads
ViT-T (encoder) 12 192 4 3
decoder 2 128 4 16
Appendix C Details of Training Hyperparameters
In Tables 9 and 10, we provide the default hyperparameters used in our experiments. In addition, we use different patch sizes for different datasets: patch size 2 for CIFAR-10, 4 for Tiny-ImageNet, and 16 for ImageNet-1K. Using smaller patch sizes increases the time consumption when calculating self-attention, but MIMIR pre-training discards 75% patches, making it still efficient.
Table 9: Pre-training hyperparameters.
Config Value
optimizer AdamW
base learning rate 1.5e-4
weight decay 0.05
optimizer momentum $\beta_{1}=0.9,\beta_{2}=0.95$
batch size 512(CIFAR-10, Tiny), 2,048 (ImageNet-1K)
learning rate schedule cosine decay
warmup epochs 40
training epochs 800
augmentation RandomResizedCrop, RandomHorizontalFlip
Table 10: Fine-tuning hyperparameters.
Config Value
optimizer AdamW
base learning rate 0.5e-2 (CIFAR-10), 1e-3 (ImageNet-1K, Tiny)
weight decay 0.05
optimizer momentum $\beta_{1}=0.9,\beta_{2}=0.999$
layer-wise lr decay 0.65, 1.0
batch size 128 (CIFAR-10), 256 (Tiny), 1,024 (ImageNet-1K)
learning rate schedule cosine decay
warmup epochs 10, 20
training epochs 100, 300
augmentation RandomResizedCrop, RandomHorizontalFlip
drop path 0.1
Table 11: Different learning rates. Fine-tuned for 50 epochs.
Dataset Models LR Natural PGD10
CIFAR-10 ViT-T 5.0e-4 76.30 47.60
1.0e-3 80.69 49.56
1.0e-2 85.62 48.78
5.0e-2 85.12 50.30
1.0e-1 84.51 50.40
Hardware. Experiments are performed on a single machine with 4 RTX A6000 (48GB) and 4 RTX A5000 (24GB) GPUs, CUDA 12.0.
@@ Appendix C Details of Training Hyperparameters
Figure 4: Natural and adversarial accuracy of ViT-S and ViT-B that are fine-tuned with different numbers of pre-train epochs under a 20-step PGD attack.
Appendix D Training Epochs Evaluation Study
Pre-training epoch. Our experiments so far are based on 800-epoch pre-training. In Figure 4, we show the influence of different numbers of epochs. We use ViT-T and ViT-S as the models for CIFAR-10, ViT-S, and ViT-B for Tiny-ImageNet. We fine-tune each model 50 epochs after MIMIR pre-training. Both adversarial and natural accuracy are improved with a longer training schedule. On CIFAR-10, the improvement of ViT-T is more significant than that of ViT-S. On Tiny-ImageNet, the increase is more obvious. Therefore, a longer pre-training schedule obviously increases performance.
Fine-tuning epoch. To show that the performance of the final model is largely attributed to MIMIR pre-training, we do a very short fine-tuning for pre-trained models. This is to train the randomly initialized classification layer since we do not have the classification layer at pre-training. Therefore, we evaluate the pre-training performance. In Figure 5, we show that MIMIR pre-training plus 5 or 10 epochs of fine-tuning is enough to achieve similar performance compared to 100-epoch fine-tuning.
Table 12: The average time consumption of a single epoch on 4 GPUs. The “mem.” refers to GPU memory usage.
CIFAR-10 Tiny-ImageNet ImageNet-1K
Architecture #Parames (M) Method time[S] mem.[GB] time[S] mem.[GB] time[S] mem.[GB]
ViT-S 21.34 PGD10 AT 149.3 2.54$\times$4 306.0 3.99$\times$4 2251.7 12.5$\times$4
FastAT 43.3 2.54$\times$4 67.7 4.03$\times$4 555.5 10.4$\times$4
MAE 16.1 3.24$\times$4 33.0 3.27$\times$4 269.6 11.1$\times$4
MIMIR 18.4 3.12$\times$4 40.0 3.18$\times$4 275.5 11.1$\times$4
ViT-B 85.27 PGD10 AT 361.2 5.39$\times$4 1022.2 8.30$\times$4 5416.7 22.1$\times$4
FastAT 122.8 5.36$\times$4 180.2 8.34$\times$4 1361.3 19.8$\times$4
MAE 53.0 5.95$\times$4 106.5 5.95$\times$4 490.9 17.0$\times$4
MIMIR 59.0 6.08$\times$4 122.0 6.11$\times$4 509.9 17.0$\times$4
ConViT-S 27.05 PGD10 AT 442.5 6.64$\times$4 897.0 12.19$\times$4 6626.5 32.5$\times$4
FastAT 106.5 5.86$\times$4 183.2 10.62$\times$4 1431.2 26.4$\times$4
MAE 33.0 10.6$\times$4 67.5 10.61$\times$4 609.7 27.5$\times$4
MIMIR 45.0 10.4$\times$4 90.0 10.54$\times$4 611.1 28.3$\times$4
Appendix E Efficiency
Table 12 provides the time consumption and memory usage of MIMIR and adversarial training at fine-tuning. The time consumption and memory usage are evaluated on 4 A6000 GPUs. Compared to 10-step PGD and FastAT, MIMIR shows much less time consumption. At pre-training, MIMIR shows slightly higher time consumption than MAE, resulting from generating adversarial examples and calculating MI between $x+\delta$ and $z$. Still, as we use a one-step algorithm, generating adversarial examples does not cost too much time. We also consider using a 10-step PGD at pre-training. Table 5 shows that using a one-step algorithm with larger $\alpha=10$ provides better performance. In addition, pre-training with MIMIR also costs less memory. Note in Table 12 that the batch size for MIMIR pre-training is much larger than the batch size for adversarial fine-tuning. In Table 12, a single epoch of MIMIR costs about $\frac{1}{8}$ time consumption compared to 10-step PGD adversarial training, and about $\frac{1}{2}$ compared to 1-step fast adversarial training (Wong, Rice, and Kolter 2020). Therefore, MIMIR is more promising regarding training costs when conducting adversarial training.
Appendix F Mutual Information and HSIC
MI measures the mutual dependence between two random variables, $X$ and $Y$. It can be decomposed as:
$I(X,Y)$ $=H(X)-H(X|Y),$ (20)
$=H(Y)-H(Y|X),$
$=H(X)+H(Y)-H(X,Y),$
where $H(X)$ and $H(Y)$ are the information entropies, $H(X|Y)$ and $H(Y|X)$ are the conditional entropies, and $H(X,Y)$ is the joint entropy of $X$ and $Y$.
Unfortunately, estimating MI in high-dimensional space is a difficult task since it may involve a precise estimation of the underlying data distribution $P_{(X,Y)}$ or $P_{(X)}$ and $P_{(Y)}$. To address this issue, the deterministic information bottleneck (DIB) (Yu, Yu, and Príncipe 2021) uses the recently proposed matrix-based Rényi’s $\alpha$-entropy functional $I_{\alpha}$ (Sanchez Giraldo, Rao, and Principe 2015; Yu et al. 2019), which suggests similar quantities to $I(X,Y)$ in terms of the normalized eigenspectrum of the Hermitian matrix of the projected data in the reproducing kernel Hilbert space (RKHS), but avoids density estimation.
Specifically, given $N$ pairs of samples $(x_{i},y_{i})_{i=1}^{N}$ (in our setup, $N$ refers to the mini-batch size), we can obtain two Gram (or kernel) matrices $K_{x}$ and $K_{y}$, for variables $X$ and $Y$, respectively, with $(K_{x})_{i,j}=\kappa_{x}(x_{i},x_{j})$, $(K_{y})_{i,j}=\kappa_{y}(y_{i},y_{j})$, in which $\kappa_{x}$ and $\kappa_{y}$ are corresponding kernel functions. The information entropy of $X$ can be expressed as:
$H_{\alpha}(X)=\frac{1}{1-\alpha}\log_{2}\left(\text{tr}(\tilde{K_{x}}^{\alpha})\right)=\frac{1}{1-\alpha}\log_{2}\left(\sum_{i=1}^{N}\lambda_{i}(\tilde{K_{x}})^{\alpha}\right),$ (21)
where $\tilde{K}$ is the normalized version of $K$, i.e., $\tilde{K}=K/{\text{tr}(K)}$, and $\lambda_{i}(\tilde{K})$ denotes the $i$-th eigenvalue of $\tilde{K}$.
Further, the joint entropy for $X$ and $Y$ can be expressed as:
$H_{\alpha}(X,Y)=H_{\alpha}\left(\frac{K_{x}\circ K_{y}}{\text{tr}(K_{x}\circ K_{y})}\right),$ (22)
where $K_{x}\circ K_{y}$ denotes the Hadamard product between the matrices $K_{x}$ and $K_{y}$.
Given Eqs. (21) and (22), the matrix-based Rényi’s $\alpha$-order mutual information $I_{\alpha}(X;Y)$ in analogy of Shannon’s MI is given by:
$I_{\alpha}(X;Y)=H_{\alpha}(X)+H_{\alpha}(Y)-H_{\alpha}(X,Y).$ (23)
Throughout this paper, we use the radial basis function (RBF) kernel $\kappa(x_{i},x_{j})=\exp(-\frac{\|x_{i}-x_{j}\|^{2}}{2\sigma^{2}})$ with kernel width $\sigma$ to obtain the Gram matrices.
The Hilbert–Schmidt Independence Criterion (HSIC) (Gretton et al. 2005) is also a kernel-based dependence measure and is usually used as a surrogate of MI. Formally, the HSIC is defined as the squared norm of the cross-covariance operator $||C_{XY}||^{2}$:
$\begin{split}&\text{HSIC}_{P_{X,Y}}(X,Y)\\ &=||C_{XY}||^{2}\\ &=\mathbb{E}_{xyx^{\prime}y^{\prime}}[\kappa_{x}(x,x^{\prime})\kappa_{y^{\prime}}(y,y^{\prime})]\\ &+\mathbb{E}_{xx^{\prime}}[\kappa_{x}(x,x^{\prime})]E_{yy^{\prime}}[\kappa_{y}(y,y^{\prime})]\\ &-2\mathbb{E}_{xy}[\mathbb{E}_{x^{\prime}}[\kappa_{x}(x,x^{\prime})]\mathbb{E}_{y^{\prime}}[\kappa_{y}(y,y^{\prime})]],\end{split}$ (24)
where $\kappa_{x}$ and $\kappa_{y}$ are kernel functions, $\mathbb{E}$ is the expectation, $x^{\prime}$ and $y^{\prime}$ are independent copies of $x$ and $y$, respectively.
Given $N$ pairs of samples $(x_{i},y_{i})_{i=1}^{N}$, the empirical estimator of HSIC is given by:
$\widehat{\text{HSIC}}_{P_{X,Y}}(X,Y)=\frac{1}{N^{2}}\text{tr}\left(K_{x}HH_{y}H\right),$ (25)
in which $(K_{x})_{i,j}=\kappa_{x}(x_{i},x_{j})$, $(K_{y})_{i,j}=\kappa_{y}(y_{i},y_{j})$, and $H=I-\frac{1}{N}\mathds{1}\mathds{1}^{T}$ is the centering matrix.

### v5

@@ IV-C Ablation Study and Further Analysis
100 $\rightarrow$ 300 epochs of FT 71.00 46.10
MI measure. In Section III-D, we provide lower and upper bound (Eq. (14)) of $I(x+\delta,z)$. According to the two bounds, $I(x+\delta,z)$ is supposed to decrease while the autoencoder learns to reconstruct the natural image $x$. This motivates us to directly embed $I(x+\delta,z)$ as a minimizing learning objective. In this paper, we use $I_{\alpha}$ [48] and HSIC [80] as estimators (detailed definitions in Appendix -J). Table VIII demonstrates the performance with different values of $\lambda$. According to the results, we use HSIC with $\lambda=1e-05$ for all other experiments.
In addition, Figure 4 provides the quantities of HSIC values while pre-training with or without using MIMIR. It is clear that MIMIR can help to decrease the mutual information between adversarial perturbation and the learned features, i.e., $I(x+\delta,z)$.
Figure 4: The HSIC values (we use HSIC as an alternative to MI) while pre-training ViT-S with and without using MIMIR.
TABLE VIII: Comparison between HSIC [51] and $I_{\alpha}$ [48] using ViT-T on CIFAR-10. Models are pre-trained 800 epochs and adversarially fine-tuned with 10-step PGD for 50 epochs.
@@
1-step is better than the 10-step of AT in pre-training. We also show that MIMIR outperforms original MAE [37] and adv MAE with different PGD steps (to generate adversarial examples for training). MAE in Table IX refers to using the original MAE for pre-training and then fine-tuning with 10-step PGD. The adv MAE refers to using adversarial examples without the MI $I(x+\delta,z)$ in the loss. The adv MAE (10-steps) refers to using the 10-step PGD algorithm ($\epsilon=8,\alpha=2$) to generate adversarial examples at pre-training. The adv MAE provides higher accuracy than MAE, which supports our statement that using adversarial examples in Masked Image Modeling creates a more difficult reconstruction task. This more difficult task further improves the performance of downstream models (see also Table XII). We use the default learning rate (i.e., 5.0e45.0e-4) of MAE, so there is a performance drop in experiments in Tables VIII and IX since AT prefers larger learning rates for CIFAR-10, as shown in Table XVII in the appendix.
TABLE IX: Comparison between different pre-training settings. All models are pre-trained for 800 epochs and then fine-tuned with 10-step PGD for 50 epochs using ViT-T on CIFAR-10.
@@
Data augmentation is not always harmful. Prior research [9, 20] has established that strong data augmentation techniques can adversely affect ViTs during adversarial training, as they may make training samples challenging to learn. However, we observe that strong data augmentation does not impair model performance when combined with longer fine-tuning periods.
The strong data augmentation refers to the combination of Randaugment [28], CutMix [29], and MixUp [30]. In this section, we evaluate two different solutions to ease this problem. First, we only use simple data augmentation for adversarial training, including random crop (or random resize crop for ImageNet-1K) and random horizontal flip (“weak aug”). Second, we use a 10-epoch warmup procedure for strong data augmentation. The warmup of Randaugment is implemented by progressively increasing the distortion magnitude from 1 to 9 (“warmup aug”). For CutMix and MixUp, we warm up by increasing the mixup probability from 0.5 to 1.0. As shown in Figure 6, “weak aug” provides the best accuracy. The “warmup aug” shows a slightly improved accuracy compared to fusing strong augmentation. Therefore, we provide a different result from [9] on the smaller dataset CIFAR-10, i.e., we show that weak augmentation is better than warmup augmentation. Even data augmentation with reduced amplitude is still difficult to learn at the beginning of adversarial training. Although strong augmentation is harmful to a normal training schedule, we show in Table X that CutMix [29], MixUp [30], and Randaugment [28] increase the accuracy of adversarial training when training with a longer schedule, e.g., 800 epochs of fine-tuning. We conjecture that combining data and strong augmentations is helpful but difficult for adversarial training to learn. Thus, more epochs are needed to learn meaningful representation. Loss and accuracy curves can be found in Appendix -E.
TABLE X: Data augmentation with longer fine-tuning schedule.
Arch Epoch Augmentation Natural PGD20
ViT-B 800 Weak Augmentation 89.90 60.26
+ CutMix [29],MixUp [30] 91.01 60.62
+ Randaugment [28] 90.19 62.75
TABLE XI: Adversarial accuracy by adaptive attacks. The models are pre-trained for 800 epochs by MIMIR and fine-tuned for 100 epochs by 1-step PGD AT.
@@
IV-D Adaptive Attacks
We evaluate MIMIR against adaptive adversaries following common practices [81]. Adaptive adversaries possess the capability to devise targeted attacks specifically tailored to exploit the mechanisms of MIMIR, particularly if they have prior knowledge of its architecture and defensive strategies. For example, the adversary may attack feature space [82, 83] since MIMIR trains the backbone to extract robust features. Here, the backbone refers to the ViT model without the classification layer, i.e., the encoder of MIMIR.
We provide two adaptive attacks specifically designed against MIMIR. First, we introduce the PGD Mutual Information attack (PGD-MI), which utilizes the MI $I(x+\delta,z)$ to generate adversarial examples, as $I(x+\delta,z)$ is used in MIMIR pre-training as a penalty in the loss. PGD-MI attacks the model by directly increasing the MI $I(x+\delta,z)$. Specifically, we add the MI loss into the PGD algorithm:
@@ IV-D Adaptive Attacks
where the value of $\lambda$ in MIMIR pre-training is available to adversaries. Second, we introduce a PGD feature attack (PGD-fea) that directly attacks the feature extracted by ViT backbones following [82]. In particular, we attack the feature extractor from the backbones after the adversarial fine-tuning. The PGD-fea attack increases the Euclidean distance between features extracted from natural and adversarial examples. We implement it using the PGD algorithm:
@@
Both PGD-MI and PGD-fea are optimized for 100 steps to ensure the attacking algorithm converges. The perturbation budget is the same as the previous evaluation, i.e., $\epsilon=8/255$ for CIFAR-10 and Tiny-ImageNet, and $\epsilon=4/255$ for ImageNet-1K.
Table XI demonstrates the adaptive evaluation results for PGD-MI and PGD-fea attacks. PGD-MI performs slightly better than the standard PGD attack, which means MI is exploitable information for perturbation crafting, but cannot significantly reduce the robustness. Furthermore, MIMIR-trained models are converged to a local optimal where the majority of predictions are constantly around the ground truth within the ball function of $\epsilon$. This also explains the resilience against both PGD and PGD-MI variants. Regarding PGD-fea, it aims to maximize feature-space divergence rather than the distance of output logits. However, MIMIR learned robust features that cannot be easily separated, so PGD-fea performs worse than PGD and PGD-MI. The collective results demonstrate that MIMIR’s IB framework induces robust learning dynamics that resist both output-space and feature-space attacks.
Figure 6: Training loss and natural accuracy of ViT-S with three different data augmentations on CIFAR-10.
Figure 7: The loss landscapes of MIMIR and MAE pre-trained models.
TABLE XII: Natural accuracy of MAE and MIMIR (800 epochs pre-training for both) that are fine-tuned on natural images.
Architecture Pre-train CIFAR-10 Tiny-ImageNet ImageNet-1K
@@
IV-E Fine-tuning with Natural Images
In Table XII, we show the results of MIMIR when fine-tuning with natural images. We compared the performance with MAE [37]. We fine-tune for 50 epochs for CIFAR-10 and Tiny-ImageNet, and 100 epochs for ImageNet-1K. The results in Table XII are reported with 800 pre-training epochs. The base learning rate used in Table XII is 0.001. The fine-tuning batch size is 512 for CIFAR-10 and Tiny-ImageNet, and 1024 for ImageNet-1K. We use weak data augmentation (“weak aug”), which includes random crop and random horizontal flip. Not surprisingly, MIMIR outperforms MAE when fine-tuning with natural data. This is because MIMIR creates a harder learning task, which is helpful to learn more discriminative representation, as we discussed in Section III-B, Design Intuition.
According to Tables XII and XIII, MIMIR consistently shows improved performance on natural data. Although the models in Table XIII show poor robustness due to fine-tuning on natural data, MIMIR pre-trained ones provide slightly better robustness. We want to clarify that poor robustness is expected when fine-tuning with natural data. First, it is known that standard training on natural data learns non-robust features [84], which hurts performance under adversarial attacks. Second, MIMIR pre-training is implemented using MSE loss plus an MI penalty between natural inputs and adversarial images. The adversarial perturbations and MI penalty help MIMIR create a more difficult and discriminative learning task to learn meaningful and robust features. This process does not include the classification layer of the final model. Therefore, MIMIR needs a fine-tuning process for superior performance on natural data and adversarial inputs. In other words, the superior performance of our experiments comes from the combination of MIMIR and the simple fine-tuning process.
IV-F Visualization of the Loss Landscape
@@
To show that the robustness of MIMIR-trained models does not stem from gradient masking, we plot the loss landscape [85] in Figure 7. The loss landscape is the visualization of the loss function as parameters change. The basic idea is to plot the loss around the optimal parameters. Formally, we consider in the 2D case,
$f_{l}(\alpha,\beta)=L(\theta^{*}+\alpha\theta_{1}+\beta\theta_{2}),$ (23)
where $\theta_{1}$ and $\theta_{2}$ are two direction vectors, $\alpha$ and $\beta$ are two arguments of $f_{l}$. In practice, we use the parameters of trained models, i.e., $\theta^{*}$. The landscapes of all models are smooth, i.e., the gradient at a certain point is clear and can also be easily estimated by local average gradients, which means the gradient is masked.
@@
TABLE XIII: Natural and adversarial accuracy of ViT-S that is fine-tuned for 5 or 50 epochs with natural images.
Fine-tune Dataset Pre-train Natural PGD
5 epochs CIFAR-10 MIMIR 93.22 0.55
MAE 93.16 0.01
Tiny-ImageNet MIMIR 63.65 0.00
MAE 62.21 0.00
ImageNet-1K MIMIR 72.45 0.20
MAE 69.47 0.02
50 epochs CIFAR-10 MIMIR 95.95 0.28
MAE 95.95 0.29
Tiny-ImageNet MIMIR 71.14 0.00
MAE 70.00 0.00
ImageNet-1K MIMIR 78.69 0.18
MAE 77.45 0.05
V Related Work
V-A Vision Transformer
The transformers [86] were first proposed in natural language processing (NLP). With the mechanism of global self-attention, transformers can effectively capture the non-local relationships among all text tokens [87, 88, 89]. A substantial effort is made to apply the transformer and self-attention mechanism in computer vision [3, 90, 5]. The pioneering work, ViT [3], demonstrated that the pure transformer architecture could achieve competitive performance on various tasks. ViT also reveals that transformers lack inductive biases [3]. For example, locality, two-dimensional neighborhood structure, and translation equivariance are inherent to CNNs but not applicable to ViTs [3]. Due to this shortcoming, ViTs usually require large-scale training to get competitive performance, such as pre-training on ImageNet-21K [36] and JFT-300M [91]. To alleviate the ViT need for large datasets, DeiT [92] introduced a teacher-student strategy to distill knowledge from a teacher CNN for a student ViT. In the MIM field, MAE [37] uses a masked autoencoder with a lightweight decoder as a visual representation learner. Its learning objective is to reconstruct the original image by the decoder while using masked images as input to the autoencoder. The advantage is that MAE can randomly discard 75% image patches when pre-training under ImageNet-1K [36], which means more efficient training.
V-B Adversarial Attacks on ViTs
The concept of adversarial attacks first appeared in [43], which proposed a formal framework and algorithms against the adversarial spam detection domain. Then, the adversarial attacks were popularized by Biggio et al. [14] and Szegedy et al. [15] in image classification. The generation of adversarial examples depends on the model’s gradient or estimated gradient in a black box situation [93]. Therefore, adversarial attacks can be easily applied to transformers by using the gradient of attention blocks concerning inputs. This also raises the question of whether transformers are more robust than CNNs. Benz et al. [94] found that CNNs are less robust than ViTs due to their shift-invariant property. Bhojanapalli et al. [95] found that ResNet models are more robust than transformers at the same model size under FGSM attack, but under PGD [21] attack, transformer models show better robustness. As ViTs process the input image as a sequence of patches, Gu et al. [13] found that ViTs are more robust than CNNs to naturally corrupted patches because the attention mechanism helps ignore naturally corrupted image patches. The later work [9] revealed that CNNs could be as robust as ViTs against adversarial attacks if CNNs are trained with proper hyperparameters.
V-C Adversarial Defense
PGD [21] adversarial training is considered one of the most effective defenses for CNNs and can withstand adaptively designed attacks [96]. However, PGD AT is harmful to the accuracy of clean data [21, 26, 2]. Generalist [1] solves this problem by formulating different training strategies for robust and natural generalization separately. DBAT [2] solves the decrease in natural accuracy by adding dummy classes [97] to the classification space.
Due to the difference between CNNs and ViTs, there have been some recent efforts to explore new adversarial training approaches for ViTs [11, 20, 12]. Mo et al. [11] presented a new adversarial training strategy based on the following observations: 1) pre-training with natural data can provide better robustness after adversarial fine-tuning, 2) gradient clipping is necessary for adversarial training, and 3) using SGD as the optimizer is better than Adam. Debenedetti et al. [20] also presented an improved training strategy for ViTs by evaluating different combinations of data augmentation policies. As adversarial training is time-consuming, AGAT [12] leverages the attention score while training to discard non-critical image patches after every layer. Unlike previous works, we provide a different training paradigm by using MIM for adversarial pre-training. Our method is efficient as we discard 75% image patches while pre-training. Our method is effective as we eliminate the information of adversarial perturbations from two information sources of natural and adversarial inputs. We also provide theoretical proof that the information of adversarial perturbations is eliminated.
V-D Self-Supervised Adversarial Pre-Training
Self-supervised learning [98, 99, 37, 100] refers to extracting meaningful representation from unlabeled data, which can be used for downstream recognition tasks. Self-supervised methods are beneficial for out-of-distribution detection on difficult, near-distribution outliers [27], which leads to using self-supervised training to improve adversarial robustness [101, 27, 102, 103, 104, 105]. The basic idea is to build a min-max learning object similar to traditional adversarial training. For example, Jiang et al. [102] considered using two adversarial samples or combining one adversarial sample and one natural sample to learn a consistent representation in contrastive learning. In more recent work, You et al. [106] proposed NIM De3 to denoise adversarial perturbations. However, the motivation of these works relies on complex self-supervised pre-training technologies, making it more difficult to understand the inner mechanisms or provide theoretical results. MIMIR not only provides better performance but also provides intuitive insights with theoretical motivation.
VI Discussion and Limitations
Following the principle of IB, we can intuitively consider a bottleneck between the encoder and decoder. As the reconstruction output is constrained by natural data $x$, the bottleneck will filter out information from adversarial perturbations $\delta$. We provide a theoretical guarantee of this bottleneck. In Eq. (8), we embed this bottleneck as a learning object to further improve the performance, which also confirms the correctness of our theoretical guarantee. With the two information sources of $x$ and $\delta$, the model is trained to learn the robust features from $x$ and forget the information of $\delta$ under the constraint of the reconstruction target.
While MIMIR shows better performance, there are still limitations. MIMIR is a pre-training method. Adversarial fine-tuning is necessary to build the final robust model. Thus, the shortcomings of traditional adversarial training cannot be completely avoided. In our experiments, we utilize the simple PGD algorithm for fine-tuning, but one can further improve MIMIR pre-trained models with more advanced approaches. In addition, MIMIR follows the design of MAE, and we also utilize the characteristic that ViTs can process variable-length inputs. Therefore, the current MIMIR cannot directly handle CNNs. While it is not trivial, we apply MIMIR to the latest CNN architecture by sparse convolution from SparK [40]. However, sparse convolution is not as efficient as dropping patch embeddings. We leave these limitations to future work.
VII Conclusions
This paper provides a novel theoretical analysis of AT for ViTs through the lens of IB. We found that constraining the MI between adversarial perturbations and their latent representations in ViT-based autoencoders, as governed by derived MI bounds, is critical for enhancing model robustness. Building upon this theoretical foundation, we propose MIMIR as a theoretically grounded pre-training method to improve adversarial robustness for ViTs. MIMIR operates by processing adversarial examples as inputs while reconstructing their natural data as targets. This approach leverages the inherent information bottleneck in autoencoder architectures to achieve two key objectives: (1) progressively eliminating perturbation-related information while (2) preserving the essential features of the original data distribution. Our extensive experimental evaluation demonstrates that MIMIR significantly outperforms existing adversarial training methods across multiple benchmark datasets, achieving SOTA results on ImageNet-1K. In addition, MIMIR is robust against unforeseen attacks and common corrupted data and can resist adaptive attacks.
Acknowledgment
The authors thank the anonymous reviewers for their constructive comments that helped improve this paper and the artifact. This work used the Dutch national e-infrastructure with the support of the SURF Cooperative using grant no. EINF-10853. This work was partially supported by the Horizon Europe programme under the project SHASAI (No. 101225866).